OpenStudy (anonymous):

x-3=x^2 LINEAR? Or NOT? Please help:)

jimthompson5910 (jim_thompson5910):

hint: graph x^2 - x + 3

OpenStudy (anonymous):

I graphed it! It's not linear. So the answer is: Not linear?

jimthompson5910 (jim_thompson5910):

yep, it's NOT a straight line, so it's NOT linear

OpenStudy (anonymous):

Thank you so much @jim_thompson5910 !!! So for future questions if you set up the equation like this/graph it you can tell?

jimthompson5910 (jim_thompson5910):

well I started with x-3=x^2 and got everything to one side x-3=x^2 -3=x^2-x 0=x^2-x+3 So that's how I knew to graph y = x^2 - x + 3

jimthompson5910 (jim_thompson5910):

and yes if you graph a completely straight line that goes on forever (straight) in both directions, then it's linear

jimthompson5910 (jim_thompson5910):

something like y = 2x+5 fits this description

OpenStudy (anonymous):

Thanks so much. You were very helpful! :)

jimthompson5910 (jim_thompson5910):

you're welcome